1. Given the data below, compute for the following: a) Forecasts for P5 to P12 using a 4-period simple moving average b) Foreplease i need asap answer2. Assuming the actual demand for P5 to P12 are as follows: PERIOD DEMAND 5,000 5,100 4,800 4,900 5,100 5,200 5,500 6,000 10
Answer #1
a 4 period simple moving average
Period Demand Forecast
1         5,000
2         4,500
3         4,800
4         5,100
5         5,500         4,850 (5000+4500+4800+5100)/4
6         5,200         4,975 (4500+4800+5100+5500)/4
7         4,200         5,150
8         4,000         5,000
9         5,800         4,725
10         5,400         4,800
11         5,000         4,850
12         6,000         5,050
b 4 month weighted moving average
Period Demand
1         5,000
2         4,500
3         4,800
4         5,100
5         5,500         4,880 (0.1*5000+0.2*4500+0.3*4800+0.4*5100)/4
6         5,200         5,140 (0.1*4500+0.2*4800+0.3*5100+0.4*5500)/5
7         4,200         5,230
8         4,000         4,850
9         5,800         4,450
10         5,400         4,880
11         5,000         5,120
12         6,000         5,180
c EXPONENTIAL SMOOTHING ALPHA= 0.3
Period Demand Forecast
4         5,100         5,000
5         5,500         5,030 (5100*0.3+5000*(1-0.3)
6         5,200         5,171 (5500*0.3+5030*(1-0.3)
7         4,200         5,180
8         4,000         4,886
9         5,800         4,620
10         5,400         4,974
11         5,000         5,102
12         6,000         5,071
LINEAR FORCASTING Yf=4639+62x Period Demand Forecast 1 5,000 2 4,500 3 4,800 5,100 5 5,500 4,949 6 5,200 5,011 7 4,200 5,073SUMMARY OUTPUT Regression Statistics Multiple R 0.36961 R Square 0.13661 Adjusted R Square 0.05028 Standard Error 588.341 ObsLupuvalu 9 Aliment Number 19 G10 X fax A B C D H I J = (F6+F7+F8+F9)/4 E F G Period Demand Forecast 1 5,000 2 4,500 3 4,800 4Clipboard Font Alignment Numbers Styles Cells x fc A B C D K L H I J (5000+4500+4800+5100)/4 (4500+4800+5100+5500)/4 =0.1*F22Clipboard Font Alignment Number G7 X for =F6*0.3+G6*(1-0.3) E A B C D F G H EXPONENTIAL SMOOTHING ALPHA= Period Demand Foreca